OpenStudy (anonymous):

Martha went shopping & bought each of her five nephews a gift, either a video costing $14.95 or a CD costing $16.88 She spent $76.88 on their gifts. How many videos & How many CD's did she buy?

OpenStudy (mertsj):

v+c=5 14.95v+16.88c=76.88 Multiply the second equation to get rid of the decimals Replace v in the second equation with 5-c from the first equation and solve. I believe it was 10 nephews last night.

OpenStudy (anonymous):

It wasn't me. This is my 1st day on this group. Please, go check original poster from last night!

OpenStudy (anonymous):

I still don't get it.

OpenStudy (mertsj):

Did you multiply the second equation by 100 to get rid of the decimals?

OpenStudy (anonymous):

(100)14.95v + (100)16.88c = (100)76.88 1495v + 1688c = 7688 Is that right?

OpenStudy (mertsj):

Yes.

OpenStudy (mertsj):

Now from the first equation, v=5-c replace v with 5-c and solve

OpenStudy (anonymous):

1495v + 1688c = 7688 1495(5-c) + 1688c =7688 7475-1495c+1688c=7688 7475-193c=7688 -7475 -7475 -193c=213 Is that right?

OpenStudy (mertsj):

7475-193c=7688 Should be 7475+193c

OpenStudy (anonymous):

I got the 7475+193c=7688 I messed up my signs 7475+193c=7688 -7475 -7475 193c=213 c=1.10

OpenStudy (mertsj):

Are you sure you copied the problem right? It should come out to be a whole number of videos and cd's. Someone posted this same problem last night and I thought it was 10 nephews. Please check the problem.

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

Her total is $76.68 not 76.88 as I posted,

OpenStudy (mertsj):

That works

OpenStudy (anonymous):

1495v + 1688c = 7668 1495(5-c) + 1688c =7668 7475-1495c+1688c=7668 7475+193c=7668 -7475 -7475 193c=193 So she bought 1 CD

OpenStudy (mertsj):

Yes and 4 videos

OpenStudy (anonymous):

Thank u

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

It makes sense. I have trouble setting up application problems.

OpenStudy (mertsj):

Everyone does in the beginning. The key is to keep practicing.

OpenStudy (anonymous):

How do we give medals? I am new. I would like to give u a medal